Calculate the ph of a solution prepared by mixing
Answer #1

Answer:

We know that,

[H+ ] of acid solution = 1×10-3 M

Volume of acid given is 2 mL, so

Number of moles of [H+ ] = M × Vol

= 1×10-3 * 2 = 2 * 10-3 mmol

Similarly,

p[OH] of the base solution = 14 -10 = 4

[OH-] of base solution = 1 ×10-4 M

Volume of basse solution = 3 mL, so

Number of moles of [OH- ] = 3×10-4 mmol

Therefore,

Excess quality of acid = 0.002 – 0.0003 = 0.0017 mmol

[H+ ] of resulting solution = = 3.4 * 10-4 M

Thus,

pH = - log (3.4 * 10-4 ) = 4 – log 3.4 = 3.47